# Minimality of free-boundary axial hyperplanes in high dimensional circular cones via calibration

### v1

@@ 3. Further notation and preliminary computations
Standard properties of the determinant guarantee that $\det g=|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}$. Define the multi-index $\beta=(1,\dots,n-1)$. We have
$\partial_{1}\wedge\dots\wedge\partial_{n}$ $=e_{\beta}\wedge e_{n}+A_{n}\,e_{\beta}\wedge e_{t}+\sum_{i=1}^{n-1}(-1)^{n-1-i}A_{i}\,e^{i}_{\beta}\wedge e_{t}\wedge e_{n}.$ (3.7)
Now the wedge products appearing in (3.7) are orthonormal each other, hence
$|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}=1+\sum_{i=1}^{n}A_{i}^{2}=1+\lambda^{2},$
and this proves (3.4).
@@
Denote by $\tilde{g}_{1}$ the $(n-1)\times(n-1)$ submatrix of $\tilde{g}$ obtained by cancelling the first row and the first column of $\tilde{g}$. By the definition of the metric on $\bigwedge^{k}T^{\ast}M$ given in (2.2) (in this case k=n2k=n-2) and the properties of the adjugate matrix (see [34, p. 73]), we have
@@ 4. Construction of the calibration
Applying the Cauchy-Schwartz inequality for wedge products (see [9, pag. 32], and note that $1$-forms are always simple), by (4), we obtain
@@ 5. Minimality of $E$
By applying Federer’s Theorem [26, Theorem 16.2], we then get

### v2

@@ 3. Further notation and preliminary computations
Standard properties of the determinant guarantee that $\det g=|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}$. Define the multi-index $\beta=(1,\dots,n)$. We have
$\partial_{1}\wedge\dots\wedge\partial_{n}$ $=e_{\beta}+\sum_{i=1}^{n}(-1)^{n-i}A_{i}\,e^{i}_{\beta}\wedge e_{t}.$ (3.7)
Now we observe that the wedge products appearing in (3.7) are orthonormal each other, hence $|\partial_{1}\wedge\dots\wedge\partial_{n}|^{2}=1+\sum_{i=1}^{n}A_{i}^{2}=1+\lambda^{2},$ and this proves (3.4).
@@
Denote by $\tilde{g}_{1}$ the $(n-1)\times(n-1)$ submatrix of $\tilde{g}$ obtained by cancelling the first row and the first column of $\tilde{g}$. By the definition of the metric on $\bigwedge^{k}T^{\ast}M$ given in (2.2) (in this case k=n2k=n-2) and the properties of the adjugate matrix (see [40, p. 73]), we have
@@ 4. Construction of the calibration
Applying the Cauchy-Schwartz inequality for wedge products (see [11, pag. 32], and note that $1$-forms are always simple), by (4), we obtain
@@ 5. Minimality of $E$
By applying Federer’s Theorem [31, Theorem 16.2], we then get